OpenStudy (anonymous):

help plz :: limit(((x+2sin(5x))/tan(3x)), x, 0)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{x + 2 \sin 5x}{\tan 3x} = \lim_{x \rightarrow 0} \frac{1 + 2 \frac{\sin 5x}{x}}{\frac{\tan 3x}{x}}=\frac{1 + 2\frac{5}{1}}{\frac{3}{1}}=\frac{11}{3}\]

OpenStudy (anonymous):

OMG thanks :) :)

OpenStudy (jack1):

@exraven ... that's some awesome funky maths you have there